# EFC++: Elastic Feature Consolidation with Prototype Re-balancing for Cold Start Exemplar-free Incremental Learning

### v2

@@ Appendix F Large-Scale Class- and Domain-Incremental Learning
In this section, we evaluate EFC++ on large-scale incremental learning benchmarks, integrating the results from Section VII-B of the main paper.
Experiments on ImageNet-1K Table VII completes the results presented in Table III of the main paper for ImageNet-1K. As already discussed in Section VII-B, EFC++ achieves remarkable results in the Cold Start scenario compared to all other competitors, while in the Warm Start setting, it achieves performance comparable to FeCAM and EFC. This can be explained by observing that the generalization gained on the large first task in a large-scale dataset reduces the necessity for plasticity when learning new tasks, thereby aligning the performance across these approaches. It must be noted that the results reported in [26] are slightly better in the Warm Start scenario in terms of average incremental accuracy compared to those reported in Table VII. This gap is likely due to the performance on the initial task, which has a significant impact on average incremental accuracy. Finally, as already observed in Section VII-B, R-DFCIL struggles in large-scale scenarios regardless of the number of classes in the first task. This is because its generative inversion network must learn to generate high-resolution images for 1,000 distinct classes.
TABLE VII: Large scale Class-incremental Experiments. We compare with the state-of-the-art on ImageNet-1K.
@@
TABLE VIII: Domain- and Class-incremental Experiments. We compare with the state-of-the-art on DN4IL, a balanced subset of DomainNet for incremental learning [32].
@@
Experiments on DN4IL (DomainNet). Table VIII extends Table 3 of the main paper for DN4IL by adding results for EwC, LwF, PASS, and SSRE, while also reporting the average incremental accuracy for all approaches. Among all the approaches, only FeCAM and SSRE achieves competitive performance, though it still falls short of the state-of-the-art. In Figure 14, we show the per-step accuracy across all encountered domains.
@@ Appendix H Ablation on EFM regularizer
In Section VIII-A of the main paper, we discuss what happens when we substitute EFM with three established regularizers: the Fisher Information Matrix (E-FIM, Eq. 3 main), Feature Distillation (FD, Eq. 5 main), and Knowledge Distillation (KD). Here, we expand that discussion by providing more details on the choice of hyperparameters We recall that the only modification to our training framework in this analysis concerns the regularization loss. The prototype update, which requires the EFM matrix, and the rebalancing of the classifier after training the backbone on the new classifier remain unchanged.
For this analysis, we validate the hyperparameters for each regularizer: $\lambda_{\text{E-FIM}}\in\{100,500,5000,10000,100000\}$, $\lambda_{\text{KD}}\in\{1,5,10,50,100\}$, and $\lambda_{\text{FD}}\in\{0.1,1,5,10,50,100\}$. Unlike EFM, where we have a theoretical explanation for the choice of hyperparameters (see Section VIII-C in the main paper), here we lack guidance on optimal values. Thus, in Table IV of the main paper we report the results obtained with the hyperparameters that perform best in each setting. Looking at the final per-step accuracy, we found that the best trade-off across settings is achieved with $\lambda_{\text{E-FIM}}=100,000$, $\lambda_{\text{KD}}=50$, and $\lambda_{\text{FD}}=1$.
@@ Appendix I Covariance drift
This section extend the discussion in Section VIII-E on the limitations of prototype rehearsal in EFC++ illustrating the results obtained on Tiny-ImageNet and on ImageNet-Subset CS. In particular, Table X and Table IX compare different update strategies of class statistics: our approach (class means estimated via EFM and fixed covariances, highlighted in cyan) against an idealized upper bound using real mean and covariances.
TABLE IX: Impact of Mean and Covariance Drift on EFC++ tested on ImageNet-Subset CS.
@@
J-A Warm Start Scenarios
@@
J-B Cold Start Scenarios
@@
TABLE X: Impact of Mean and Covariance Drift on EFC++ tested on Tiny-ImageNet CS.

### v3

@@ Appendix F Large-Scale Class- and Domain-Incremental Learning
In this section, we evaluate EFC++ on large-scale incremental learning benchmarks, integrating the results from Section 7.2 of the main paper.
Experiments on ImageNet-1K Table 7 completes the results presented in Table 3 of the main paper for ImageNet-1K. As already discussed in Section 7.2, EFC++ achieves remarkable results in the Cold Start scenario compared to all other competitors, while in the Warm Start setting, it achieves performance comparable to FeCAM and EFC. This can be explained by observing that the generalization gained on the large first task in a large-scale dataset reduces the necessity for plasticity when learning new tasks, thereby aligning the performance across these approaches. It must be noted that the results reported in fetril are slightly better in the Warm Start scenario in terms of average incremental accuracy compared to those reported in Table 7. This gap is likely due to the performance on the initial task, which has a significant impact on average incremental accuracy. Finally, as already observed in Section 7.2, R-DFCIL struggles in large-scale scenarios regardless of the number of classes in the first task. This is because its generative inversion network must learn to generate high-resolution images for 1,000 distinct classes.
Table 7: Large scale Class-incremental Experiments. We compare with the state-of-the-art on ImageNet-1K.
@@
Table 8: Domain- and Class-incremental Experiments. We compare with the state-of-the-art on DN4IL, a balanced subset of DomainNet for incremental learning gowda2023dual .
@@
Experiments on DN4IL (DomainNet). Table 8 extends Table 3 of the main paper for DN4IL by adding results for EwC, LwF, PASS, and SSRE, while also reporting the average incremental accuracy for all approaches. Among all the approaches, only FeCAM and SSRE achieves competitive performance, though it still falls short of the state-of-the-art. In Figure 14, we show the per-step accuracy across all encountered domains.
@@ Appendix H Ablation on EFM regularizer
In Section 8.1 of the main paper, we discuss what happens when we substitute EFM with three established regularizers: the Fisher Information Matrix (E-FIM, Eq. 3 main), Feature Distillation (FD, Eq. 5 main), and Knowledge Distillation (KD). Here, we expand that discussion by providing more details on the choice of hyperparameters We recall that the only modification to our training framework in this analysis concerns the regularization loss. The prototype update, which requires the EFM matrix, and the rebalancing of the classifier after training the backbone on the new classifier remain unchanged.
For this analysis, we validate the hyperparameters for each regularizer: $\lambda_{\text{E-FIM}}\in\{100,500,5000,10000,100000\}$, $\lambda_{\text{KD}}\in\{1,5,10,50,100\}$, and $\lambda_{\text{FD}}\in\{0.1,1,5,10,50,100\}$. Unlike EFM, where we have a theoretical explanation for the choice of hyperparameters (see Section 8.3 in the main paper), here we lack guidance on optimal values. Thus, in Table 4 of the main paper we report the results obtained with the hyperparameters that perform best in each setting. Looking at the final per-step accuracy, we found that the best trade-off across settings is achieved with $\lambda_{\text{E-FIM}}=100,000$, $\lambda_{\text{KD}}=50$, and $\lambda_{\text{FD}}=1$.
@@ Appendix I Covariance drift
This section extends the discussion in Section 8.5 on the limitations of prototype rehearsal in EFC++ illustrating the results obtained on Tiny-ImageNet and on ImageNet-Subset CS. In particular, Table 10 and Table 9 compare different update strategies of class statistics: our approach (class means estimated via EFM and fixed covariances, highlighted in cyan) against an idealized upper bound using real mean and covariances.
Table 9: Impact of Mean and Covariance Drift on EFC++ tested on ImageNet-Subset CS.
@@
J.1 Warm Start Scenarios
@@
J.2 Cold Start Scenarios
@@
Table 10: Impact of Mean and Covariance Drift on EFC++ tested on Tiny-ImageNet CS.